# Common and Rare Fundus Diseases Identification Using Vision-Language Foundation Model with Knowledge of Over 400 Diseases

### v1

@@ 3 Discussion
In this study, we trained a vision-language-foundation model RetiZero for retinal imaging using vast fundus images paired with text description. It has demonstrated its strong capability in representing retinal disease features across a wide range of downstream tasks of retinal disease identification, including internal domain and cross-domain classification, few-shot fine-tuning, zero-shot recognition, and image-to-image retrieval. The performance of RetiZero is superior to two state-of-the-art LFMs, RETFound [15] and FLAIR [16]. These results collectively demonstrated the superior generalizable and robust performance of RetiZero in both common and rare retinal disease identification.
The superiority of RetiZero over RETFound [14] and FLAIR [16] can be attributed to its unique design and diverse data used for pre-training. Although the RETFound model [14], pre-trained on a large number of fundus images using the MAE architecture, can enhance the performance of various downstream tasks, it included a limited number of fundus disease categories, particularly rare fundus diseases. In addition, it lacks the incorporation of textual information, resulting in insufficient characterization of image feature attributes, making it unsuitable for text prompt-based zero-shot fundus disease screening tasks and limiting its application in clinical practice scenarios, especially for the identification of rare fundus diseases. In contrast, FLAIR [16], based on the CLIP architecture, incorporates textual description information during network training to enhance the representation of image feature attributes. However, it is pre-trained on a very limited dataset of fundus disease knowledge, leading to poor performance in zero-shot recognition tasks for rare fundus diseases. Furthermore, FLAIR lacks guidance for learning information such as lesion contours and topological structures in images, resulting in low performance in fundus disease identification through image-to-image retrieval.
@@
Classification of retinal photographs for fundus disease identification is a well-studied task. Driven by comprehensive ophthalmic knowledge and contextual information from fundus images, RetiZero provides strong feature representation capability and robustness, enabling its superior performance for fundus disease identification across internal domain, cross-domain, and few-shot learning. Although the foundation model can reduce the sample size needed for training, classification tasks still require a certain number of images for fine-tuning and testing. It would be very challenging to collect enough sample sizes for rare diseases. To address this issue, we introduced image-to-image retrieval and zero-shot recognition tasks. Both tasks do not require data for fine-tuning, making them particularly useful for rare diseases. Image-to-image retrieval involves determining the category of a query fundus image based on feature similarity scores between the query and candidate images. RetiZero leverages its excellent image content representation capabilities of the MAE architecture and the textual feature alignment characteristics of the CLIP architecture. Therefore, it achieved superior performance in fundus disease retrieval tasks based on retinal image content. While the zero-shot learning setting allows for the recognition of rare diseases with as few as one sample. RetiZero learned textual knowledge of over 400 types of fundus diseases, enabling it to perform promisingly in zero-shot fundus disease recognition tasks, a feat not achievable by RETFound.
@@
We also recognized limitations and the need for improvements in the current work. Although our collected dataset includes knowledge about over 400 types of fundus diseases, the imbalance across different categories may limit RetiZero’s performance in downstream tasks. Therefore, further enriching the dataset with varied quantities of different categories of fundus diseases, especially for rare fundus diseases, will be part of our future work. In addition, while RetiZero has shown promising performance across multiple tasks and datasets, specialized models optimized for specific tasks may outperform generic models. Therefore, we will further explore improvements of RetiZero for specific tasks.
@@
In conclusion, the proposed feature-calibrated retinal vision-language foundation model (RetiZero) with knowledge of over 400 retinal diseases can effectively represent the rich contextual feature information in fundus images, as well effectively learn the alignment between retinal image features and textual descriptions. RetiZero achieved superior performance on feature representation and generalizability across different retinal disease recognition tasks at various ophthalmic centers, different degrees of domain drift, and very limited training samples. Particularly, the outstanding performance of RetiZero in zero-shot fundus disease identification and image-to-image retrieval-based fundus disease recognition holds significant importance for screening fundus diseases in clinical practice, especially for rare fundus diseases.
@@ 4.1.1 Data for pretraining:
We utilized RETFound [14], pre-trained on over 900,000 fundus images using the MAE architecture, as the pre-trained backbone for the Image Encoder of RetiZero. Meanwhile, we introduced low-rank learnable factors into the pre-trained RETFound and leveraged the CLIP architecture to learn image-text knowledge, aiming to enhance the model’s understanding of image-text correlations and improve its feature representation capabilities. We pre-trained RetiZero using our collected dataset comprising 341,896 image-text pairs, covering over 400 fundus diseases. Since the dataset used for pre-training with the MAE architecture has been previously described in RETFound [14], this paper focuses on detailing the 341,896 image-text dataset we collected. As shown in Supplementary Table 13, the image-text pretraining data mainly consists of three parts: publicly available dataset with category information, data from the ophthalmic books with description information, and data from online resources with description. Specifically, we collected a total of 303,129 fundus images from 29 publicly available datasets, covering over 100 different categories of retinal diseases. We used these category labels as textual descriptions corresponding to the fundus images input into RetiZero. To enable RetiZero to acquire a more comprehensive knowledge of ophthalmology, we invited 10 ophthalmologists to further collect 23,228 fundus images with corresponding textual descriptions from 180 ophthalmic books. As shown in Supplementary Table 14, these images cover 414 ophthalmic descriptive labels, encompassing nearly all known fundus diseases to date. Furthermore, we also collected 28,800 fundus data with relevant descriptions from the online resources. We assembled a team of 12 ophthalmologists to manually clean and organize 15,544 images along with their corresponding textual descriptions. In summary, the dataset for pretraining RetiZero covers almost all currently known fundus diseases, integrating very comprehensive ophthalmic knowledge. We pre-trained RetiZero on the public platform Pytorch and Nvidia Geforce DGX A100 GPU (80G). The batch size was set to 128. Adam was adopted as the optimizer to optimize RetiZero. The procedure of data collection for RetiZero pretraining was provided in Figure 1.
4.1.2 Data for internal domain retinal diseases identification:
@@
To verify the performance of the proposed RetiZero in the task of retinal disease identification, we built three datasets across multiple ophthalmic centers: health dataset 1 (H1), health dataset 2 (H2), and health dataset 3 (H3). This study was approved by the Joint Shantou International Eye Center Institutional Review Board and adhered to the principles of the Declaration of Helsinki. The data has been de-identified. In accordance with IRB regulations, if the data does not contain any identifiable patient information, informed consent is not required. As a result, this study has been granted approval to waive the need for informed consent. The clinical assessment and labeling procedure are shown in Supplementary Figure 11. H1 dataset, consists 11,414 fundus images covering 14 categories of retinal diseases and a normal condition, was collected from multiple eye clinics using different fundus cameras. We further divided H1 into training (6,942), validation (2,284), and testing (2,288) for model fine-tuning, model selection, and performance verification, respectively. More details of data information are given in Supplementary Table 4. H2 dataset consists of 7,812 fundus images, including 12 types of retinal diseases and 1 normal condition. The category and data information are given in Supplementary Table 5. To validate the performance of fine-tuning RetiZero for retinal disease identification on the H2 dataset, we partitioned the H2 dataset into training (4,682), validation (1,561), and testing (1,569) sets, respectively, for model fine-tuning, model selection, and performance evaluation. Supplementary Table 6 provides the category and data distribution information for the H3 dataset, which comprises 10,863 fundus images across 12 categories. We divided the H3 dataset into training (6,511), validation (2,174), and testing (2,178) sets for model fine-tuning, selection, and performance evaluation. In this paper, we fine-tuned RetiZero to the task of internal domain retinal disease identification on the public platform Pytorch and Nvidia Geforce 3090 GPU (24). Adam optimizer and cross-entropy loss function were adopted to guide the model fine-tuning. The total iteration epoch and batch size were set to 100 and 64, respectively.
@@ 4.1.3 Data for few-shot fine-tuning:
To evaluate the performance of RetiZero in the few-shot fine-tuning downstream task, we further reorganized the H1, H2, and H3 datasets. Specifically, we randomly selected 5 samples from each category of H1 training set, H2 training set, and H3 training set for few-shot fine-tuning, while retaining the validation and testing datasets for model selection and performance evaluation. More details about category and data distribution information are given in Supplementary Table 7 to Supplementary Table 9. In this experiment, RetiZero was fine-tuned on the public platform Pytorch and Nvidia Geforce 3090 GPU (24G). Adam optimizer and cross-entropy loss function were adopted to guide the model optimization. The total iteration epoch and batch size were set to 1000 and 32, respectively.
@@ 4.1.4 Data for cross-domain fundus disease identification:
We combined three datasets from different hospitals, H1, H2, and H3, into a dataset named EYE-15, containing 30,089 fundus images that include 14 common fundus diseases and 1 normal category. This dataset was used to validate RetiZero’s performance in screening common fundus diseases in zero-shot and image-to-image retrieval approaches. The data distribution of each category in EYE-15 was provided in Supplementary Table 1. Moreover, we further collaborated with several ophthalmologists from multiple eye clinics to collect 7,007 fundus images by different fundus cameras (EYE-52 dataset), comprising 51 fundus diseases and 1 normal condition, to validate the performance of zero-shot fundus disease recognition and fundus disease identification by image-to-image retrieval in a more challenging setting. The data distribution was shown in Supplementary Table 2. As shown in Supplementary Table 2, EYE-52 comprises many clinically particularly rare fundus diseases, such as albinism, Bietti crystalline dystrophy, choroidal coloboma, and choroidal neoplasm. We adopted Top1, Top3, and Top 5 accuracy to evaluate the performance of RetiZero in both tasks of zero-shot fundus disease recognition and fundus disease identification by image-to-image retrieval. And, Supplementary Figure 11 provides the process of the collection for the EYE-15 and EYE-52 datasets. In this paper, we also adopted Precsion@1, Precesion@3, and Precsion@5 as the metrics to evaluate the performance of different foundation models in the task of fundus disease retrieval. Precision@N is a metric used to evaluate the performance of information retrieval systems and ranking algorithms. It is specifically used to measure the precision of the top $N$ results returned by a system. Here is the formula and its explanation:

### v2

@@ 3 Discussion
In this study, we trained a vision-language-foundation model RetiZero for retinal imaging using vast fundus images paired with text description. Comprehensive experimental results demonstrated that RetiZero has a strong capability in representing retinal disease features across a wide range of downstream tasks of retinal disease identification, including internal domain and cross-domain classification, few-shot fine-tuning, zero-shot recognition, and image-to-image retrieval. The performance of RetiZero is superior to two state-of-the-art ophthalmic LFMs, RETFound [15] and FLAIR [17]. These results collectively demonstrated the superior generalizable and robust performance of RetiZero in both common and rare retinal disease identification.
The superiority of RetiZero over RETFound [15] and FLAIR [17] can be attributed to its unique design and diverse data used for pre-training. Although the RETFound model [15], pre-trained on a large number of fundus images using the MAE architecture, can enhance the performance of various downstream tasks, it includes a limited number of fundus disease categories, particularly rare fundus diseases. In addition, it lacks the incorporation of textual information, resulting in inadequate characterization of image feature attributes, making it unsuitable for text prompt-based zero-shot fundus disease screening tasks and limiting its application in clinical practice scenarios, especially for the identification of rare fundus diseases. In contrast, FLAIR [17], based on the CLIP architecture, incorporates textual description information during network training to enhance the representation of image feature attributes. However, it is pre-trained on a very limited dataset of fundus disease knowledge, leading to poor performance in zero-shot recognition tasks for rare fundus diseases. Furthermore, FLAIR lacks guidance for learning information such as lesion contours and topological structures in images, resulting in low performance in fundus disease identification through image-to-image retrieval.
@@
Classification of retinal photographs for fundus disease identification is a well-studied task. Driven by comprehensive ophthalmic knowledge and contextual information from fundus images, RetiZero provides strong feature representation capability and robustness, enabling its superior performance for fundus disease identification across internal domain, cross-domain, and few-shot learning. Although the foundation model can reduce the sample size needed for training, classification tasks still require a certain number of images for fine-tuning and testing. It would be very challenging to collect enough sample sizes for rare diseases. To address this issue, we introduced image-to-image retrieval and zero-shot recognition tasks. Both tasks do not require data for fine-tuning, making them particularly useful for rare diseases. Image-to-image retrieval involves determining the category of a query fundus image based on feature similarity scores between the query and candidate images. RetiZero leverages its excellent image content representation capabilities of the MAE architecture and the textual feature alignment characteristics of the CLIP architecture. Therefore, it achieved superior performance in fundus disease retrieval tasks based on retinal image content. While the zero-shot learning setting allows for the recognition of rare diseases with even as few as one sample. RetiZero learned textual knowledge of over 400 types of fundus diseases, enabling it to perform promisingly in zero-shot fundus disease recognition tasks, a feat not achievable by RETFound. By analyzing the two rounds of clinical evaluations conducted by 19 ophthalmologists from Singapore, the United States, and China, we observed that RetiZero’s zero-shot Top3 and Top5 accuracies were comparable to and exceeded those of most ophthalmologists, respectively. Particularly, RetiZero’s image-to-image retrieval-based Top1 accuracy outperformed that of most ophthalmologists. More interestingly, after introducing RetiZero’s top five predictions as an aid, 18 out of 19 ophthalmologists showed improved diagnostic accuracy, resulting in an overall accuracy increase of 7.6%. This highlights the significant clinical importance of RetiZero in the assistance of enhancing the diagnostic accuracy of clinicians. Furthermore, the clinical evaluation results further confirm the effectiveness of the reference answers provided by RetiZero, with the average diagnostic accuracy significantly improving from 55.2% to 62.8%, particularly benefiting less experienced ophthalmologists. Additionally, with RetiZero’s assistance, the confidence of doctors in their diagnoses also increased. Moreover, Spearman correlation analysis further demonstrated a correlation between RetiZero’s correct diagnosis ranking and the diagnostic accuracy of ophthalmologists, underscoring the huge potential of RetiZero as a valuable tool in clinical practice.
@@
We also recognized limitations and the need for improvements in the current work. Although our collected dataset includes knowledge of over 400 types of fundus diseases, the imbalance across different categories may limit RetiZero’s performance in downstream tasks. Therefore, further enriching the dataset with varied quantities of different categories of fundus diseases, especially for rare fundus diseases, will be part of our future work. In addition, while RetiZero has shown promising performance across multiple tasks and datasets, specialized models optimized for specific tasks may outperform generic models. Therefore, we will further explore improvements of RetiZero for specific tasks.
@@
In conclusion, the proposed feature-calibrated retinal vision-language foundation model (RetiZero) with knowledge of over 400 retinal diseases can effectively represent the rich contextual feature information in fundus images, as well effectively learn the alignment between retinal image features and textual descriptions. RetiZero achieved superior performance on feature representation and generalizability across different retinal disease recognition tasks at various ophthalmic centers, different degrees of domain drift, and very limited training samples. In particular, the excellent performance of RetiZero in zero-shot fundus disease identification and image-to-image retrieval-based fundus disease recognition holds significant importance for screening fundus diseases in clinical practice, especially for rare fundus diseases. Furthermore, comprehensive clinical evaluation results further demonstrated that RetiZero can assist in improving ophthalmologists’ diagnostic accuracy and confidence, particularly benefiting less experienced clinicians.
@@ 4.1.1 Data for pretraining:
We utilized RETFound [15], pre-trained on over 900,000 fundus images using the MAE architecture, as the pre-trained backbone for the Image Encoder of RetiZero. Meanwhile, we introduced low-rank learnable factors into the pre-trained RETFound and leveraged the CLIP architecture to learn image-text knowledge, aiming to enhance the model’s understanding of image-text correlations and improve its feature representation capabilities. We pre-trained RetiZero using our collected dataset comprising 341,896 image-text pairs and covering over 400 fundus diseases. Since the dataset used for pre-training with the MAE architecture has been previously described in RETFound [15], this paper focuses on detailing the 341,896 image-text datasets that we have collected. As shown in Supplementary Table 13, the image-text pretraining data mainly consists of three parts: publicly available dataset with category information, data from the ophthalmic books with description information, and data from online resources with descriptions. Specifically, we collected a total of 303,129 fundus images from 29 publicly available datasets, covering over 100 different categories of retinal diseases. We used these category labels as textual descriptions corresponding to the fundus images input into RetiZero. To enable RetiZero to acquire a more comprehensive knowledge of ophthalmology, we invited 10 ophthalmologists to further collect 23,228 fundus images with corresponding textual descriptions from 180 ophthalmic books. As shown in Supplementary Table 14, these images cover 414 ophthalmic descriptive labels, encompassing nearly all known fundus diseases to date. Furthermore, we also collected 28,800 fundus data with relevant descriptions from the online resources. We assembled a team of 12 ophthalmologists to manually clean and organize 15,544 images along with their corresponding textual descriptions. In summary, the dataset for pretraining RetiZero covers almost all currently known fundus diseases, integrating very comprehensive ophthalmic knowledge. We pre-trained RetiZero on the public platform PyTorch and Nvidia Geforce DGX A100 GPU (80G). The batch size was set to 128. Adam was adopted as the optimizer to optimize RetiZero. The procedure of data collection for RetiZero pretraining is provided in Figure 1.
4.1.2 Data for internal domain retinal disease identification:
@@
To verify the performance of the proposed RetiZero in the task of retinal disease identification, we built three datasets across multiple ophthalmic centers: health dataset 1 (H1), health dataset 2 (H2), and health dataset 3 (H3). The clinical assessment and labeling procedure are shown in Supplementary Figure 11. H1 dataset, consisting of 11,414 fundus images and covering 15 categories of retinal diseases and normal condition, was collected from different clinics using different fundus cameras. We further divided H1 into training (6,942), validation (2,284), and testing (2,288) for model fine-tuning, model selection, and performance verification, respectively. More details of data information are given in Supplementary Table 4. The H2 dataset consists of 7,812 fundus images acquired from different clinics and devices, including 12 types of retinal diseases and 1 normal condition. The category and data information are given in Supplementary Table 5. To validate the performance of fine-tuning RetiZero for retinal disease identification on the H2 dataset, we partitioned the H2 dataset into training (4,682), validation (1,561), and testing (1,569) sets, respectively, for model fine-tuning, model selection, and performance evaluation. Supplementary Table 6 provides the category and data distribution information for the H3 dataset, which comprises 10,863 fundus images across 12 categories. In general, as shown in Supplementary Figure 11, the distribution characteristics of the H1 and H2 datasets are cross-center and cross-device, whereas, in contrast to the H1 and H2 datasets, the H3 dataset is collected from a different device. We divided the H3 dataset into training (6,511), validation (2,174), and testing (2,178) sets for model fine-tuning, selection, and performance evaluation. This study was approved by the Joint Shantou International Eye Center Institutional Review Board and adhered to the principles of the Declaration of Helsinki. The data has been de-identified. In accordance with IRB regulations, if the data does not contain any identifiable patient information, informed consent is not required. As a result, this study has been granted approval to waive the need for informed consent. In this paper, we fine-tuned RetiZero to the task of internal domain retinal disease identification on the public platform PyTorch and Nvidia Geforce 3090 GPU (24). Adam optimizer and cross-entropy loss function were adopted to guide the model fine-tuning. The total iteration epoch and batch size were set to 100 and 64, respectively.
@@ 4.1.3 Data for few-shot fine-tuning:
To evaluate the performance of RetiZero in the few-shot fine-tuning downstream task, we further reorganized the H1, H2, and H3 datasets. Specifically, we randomly selected 5 samples from each category of H1 training set, H2 training set, and H3 training set for few-shot fine-tuning, while retaining the validation and testing datasets for model selection and performance evaluation. More details about category and data distribution information are given in Supplementary Tables 7 to 9. In this experiment, RetiZero was fine-tuned on the public platform PyTorch and Nvidia Geforce 3090 GPU (24G). Adam optimizer and cross-entropy loss function were adopted to guide the model optimization. The total iteration epoch and batch size were set to 1000 and 32, respectively.
@@ 4.1.4 Data for cross-domain fundus disease identification:
We combined three datasets from different hospitals, H1, H2, and H3, into a dataset named EYE-15, containing 30,089 fundus images that include 14 common fundus diseases and 1 normal category. This dataset was used to validate RetiZero’s performance in screening common fundus diseases in zero-shot and image-to-image retrieval approaches. The data distribution of each category in EYE-15 was provided in Supplementary Table 1. We combined three datasets from different hospitals, H1, H2, and H3, into a dataset named EYE-15, containing 30,089 fundus images that included 14 common fundus diseases and 1 normal category. This dataset was used to validate RetiZero’s performance in screening common fundus diseases in zero-shot and image-to-image retrieval approaches. The data distribution of each category in EYE-15 was provided in Supplementary Table 1. Moreover, we further collaborated with several ophthalmologists from multiple eye clinics to collect 7,007 fundus images by different fundus cameras (EYE-52 dataset), comprising 51 fundus diseases and 1 normal condition, to validate the performance of zero-shot fundus disease recognition and fundus disease identification by image-to-image retrieval in a more challenging setting. The data distribution was shown in Supplementary Table 2. As shown in Supplementary Table 2, EYE-52 comprises many clinically particularly rare fundus diseases, such as albinism, Bietti crystalline dystrophy, choroidal coloboma, and choroidal neoplasm. We adopted Top1, Top3, and Top 5 accuracy to evaluate the performance of RetiZero in both tasks of zero-shot fundus disease recognition and fundus disease identification by image-to-image retrieval. TopK accuracy is a metric used to evaluate the performance of a classification model by determining whether the correct label for a given input appears within the top $K$ predictions made by the model, as follows:
$TopK=\frac{Number\ of\ correct\ labels\ in\ topK\ predictions}{Total\ number\ of\ inputs},$ (1)
Supplementary Figure 11 provides the process of the collection for the EYE-15 and EYE-52 datasets. In this paper, we also adopted Precsion@1, Precesion@3, and Precsion@5 as the metrics to evaluate the performance of different foundation models in the task of fundus disease retrieval. Precision@N is a metric used to evaluate the performance of information retrieval systems and ranking algorithms. It is specifically used to measure the precision of the top $N$ results returned by a system. Here is the formula and its explanation:
@@ 4.4 Definition of Dirichlet distribution
5 Acknowledgements
This work is supported by the A*STAR Central Research Fund (“Robust and Trustworthy AI system for Multi-modality Healthcare” to H.F.), Career Development Fund (C222812010 to H.F.), and Advanced Manufacturing and Engineering (AME) Programmatic Fund (A20H4b0141 to C.Y.C.), the National Medical Research Council, Singapore (MOH-CSASI22jul-0001 to C.Y.C.), Shantou Science and Technology Program (190917085269835 to H. C.), 2020 Li Ka Shing Foundation Cross-Disciplinary Research Grant (2020LKSFG14B to H. C.), National Key R&D Program of China (2018YFA0701700 to H. C. and X. C.), the National Nature Science Foundation of China (U20A20170 to X.C.).